# AddAdditional References

### v1

@@ I Introduction
We expect these results to be relevant in many control applications where specifications are provided in the form of social cost/welfare function [13]. The mathematical guarantees that we provide can be incorporated in the design of system-wide metrics that are now being used in multi-agent control, like Price of Anarchy [14, 15, 16, 17, 18]. They can also guide the design of optimization dynamics [19, 20, 21] and resource allocation mechanisms [22, 23]. Finally, our work responds to the growing interest in fairness in control [1, 24, 25, 26, 27, 28], as many fairness notions arise naturally from careful considerations of the level of interpersonal comparability.
@@
The remainder of this paper is structured as follows. In Section II, we introduce desirable axioms that will set the stage for a principled welfarist approach based on social choice theory [32, 33, 34]. In Section III, we derive results on how desirable axioms, together with limits/possibilities of interpersonal cost comparability, determine which social cost aggregations are permissible. In Section IV, we illustrate the versatility of the proposed framework via examples from water distribution and transportation.
@@ II Welfarism
In this section we introduce preliminaries adopted from social choice theory that allow a rigorous aggregation of individual costs and derivation of social cost functions. We lay out the foundations of a so called ‘welfarist’ approach, that requires that all the relevant information for the social decision is contained in the agents’ cost functions.
@@ II-A Preliminaries and Axiomatic Foundations
[…]
$x\succsim_{i}y\quad\Longleftrightarrow\quad J_{i}(x)\leq J_{i}(y).$
If $J_{i}$ additionally satisfies appropriate continuity conditions, we say each agent’s preferences are numerically representable or cardinalizable. We let $\mathbf{J}=(J_{1},\dots,J_{n})$ denote a profile of such costs for the entire set of agents. When formulating a resource allocation problem that depends on agents’ individual evaluations represented by $\mathbf{J}$, one needs to define a single social preference relation $\succsim_{\mathbf{J}}$ on $\mathcal{X}$, so as to capture the collective or “social” viewpoint. We want to define a mapping between cost profiles and the resulting social preference, which can be expressed using Social Cost Functionals (SCFL):
Definition 1.
Let $\mathcal{R}$ be the set of all complete and transitive binary relations on $\mathcal{X}$. A Social Cost Functional (SCFL) is a mapping
$\mathfrak{F}\;:\;\mathcal{J}^{n}\;\longrightarrow\;\mathcal{R},$
such that, for each profile $\mathbf{J}=(J_{1},\dots,J_{n})\in\mathcal{J}^{n}$, it assigns a social preference relation $\succsim_{\mathbf{J}}=\mathfrak{F}(\mathbf{J})$ on $\mathcal{X}$.
@@
In control applications, it is convenient to work with real-valued measures of social cost rather than a social ordering of outcomes. This requires an additional mild continuity assumption on the SCFL, which postulates that if an outcome $x$ is socially strictly preferred over $y$, it should remain so under a small enough perturbation of the individual costs. As we show in Section III-B, this condition is easily verified for the setting considered in this work.
@@ II-B Welfarism
As mentioned in the introduction, several well-known examples of SCFs $C$ have been proposed to aggregate individual costs. For instance, the utilitarian approach, often attributed to Bentham [36] as later formalized by Harsanyi [37], aggregates costs via a sum: $C_{\mathrm{util}}(J_{1},\dots,J_{n})=\sum_{i=1}^{n}J_{i}$. By contrast, the Rawlsian or min-max rule, inspired by Rawls [38], focuses on the well-being of the worst-off agent: $C_{\mathrm{rawls}}=\max_{i\in\mathcal{N}}J_{i}$. Another example is the Nash Social Welfare function [39, 40], which is defined as $C_{\mathrm{Nash}}(u_{1},\dots,u_{n})=-\prod_{i=1}^{n}(-J_{i})$. Typically, selecting a SCF as the objective in a control problem is driven by the desired properties, such as efficiency, fairness, or equity, that the designer wishes to achieve. The focus of this work is to provide a principled way to make this choice based on the fundamental assumptions of interpersonal comparability of the agents’ costs, from which the appropriate aggregation rule naturally follows.
@@ III-A Comparability Levels
Let $C:\mathcal{X}\to\mathbb{R}$ be a social cost function defined on a set of outcomes $\mathcal{X}$. A vector of transformations $\{\phi_{i}\in\Phi:i\in\mathcal{N}\}$ is called an invariance transformation under $C$ if, for any cost profile $\mathbf{J}=(J_{1},\dots,J_{n})$, the transformed profile
@@ III-B Permissible Social Cost Functions
Let $\mathfrak{F}:\mathcal{J}^{n}\to\mathcal{R}$ be a Social Cost Functional satisfying (IIA) $+$ (P) (unless stated otherwise). We have the following results, for each level of interpersonal comparability.
@@
[…]
@@
$-\prod_{i}\left[J_{i}(x_{0})-J_{i}(x)\right]^{c_{i}}$ with benchmark $x_{0}$
$-\prod_{i}\left[-J_{i}(x)\right]^{c_{i}}$ with $b_{i}=0$ and negative costs
$\sum_{i}c_{i}\,J_{i}(x)$ $\frac{1}{n}\sum_{i}J_{i}(x)\;+\;g\left(\left[\begin{smallmatrix}J_{1}(x)-\frac{1}{n}\sum_{i}{J}(x)\\ \vdots\\ J_{n}(x)-\frac{1}{n}\sum_{i}{J}(x)\end{smallmatrix}\right]\right)$ $g$ homogeneous of degree 1
@@ III-C Admissible Operations on Social Cost Functions
One needs to be particularly careful about using SCFs for these purposes because Lemma 1 only certifies their use for selecting the best social choice. It is allowed, however, to perform other operations on the values of SCF for different outcomes, as long as the result of such operations is invariant with respect to the transformations listed in Section III-A, which guarantees that the result of such operations is meaningful. The following lemma allows us to characterize the set of operations that satisfy this property in Proposition 1.
Lemma 3.
Consider $n$ values $(w_{1},\dots,w_{n})\in\mathbb{R}^{n}$. Suppose $q(w_{1},\dots,w_{n})$ is a rational function satisfying $\forall\,a\in\mathbb{R}_{++},\;\forall\,b\in\mathbb{R}$
$q(w_{1},\dots,w_{n})\;=\;q(a\,w_{1}+b,\;\dots,\;a\,w_{n}+b)$
Then $q$ is a rational function of the ratios of differences:
$\frac{\,w_{i}-w_{j}\,}{\,w_{k}-w_{\ell}\,},\quad 1\leq i,j,k,\ell\leq n,\quad(w_{k}-w_{\ell}\neq 0).$
Conversely, any rational expression built from these ratios of differences is invariant under $t\mapsto a\,t+b$.
Proof.
[…]
$K\bigl{(}x_{1},y_{1},\dots,x_{n},y_{n}\bigr{)}^{\mathrm{GL}_{2}}\;=\;K\!\Bigl{(}\,\tfrac{f_{i,j}}{f_{k,\ell}}\Bigr{)}.$
Substituting pair of variables for the embedded representation $(x_{i},y_{i})=(w_{i},1)$ into $f_{i,j}$, i.e. expressing them as $f_{i,j}=x_{i}\,y_{j}-y_{i}\,x_{j}=w_{i}-w_{j}$, together with the fact that $\mathrm{Aff}(1,\mathbb{R})$ is a subgroup of $GL_{2}(\mathbb{R})$ gives us the result:
$K(w_{1},\dots,w_{n})^{\mathrm{Aff}(1,K)}\;=\;K\Bigl{(}\,\tfrac{\,w_{i}-w_{j}\,}{\,w_{k}-w_{\ell}\,}\Bigr{)}.\qed$
@@
We show that admissible transformations described in Section III-A induce an affine group action on the SCF $C$. For (CUC) we obtain $C(\phi_{\text{CUC}}(\cdot))=aC(\cdot)+\sum_{i}c_{i}b_{i}$. For (CFC) we obtain $C(\phi_{\text{CFC}}(\cdot))=aC(\cdot)+nb$. For (CNC) we obtain $C(\phi_{\text{CNC}}(\cdot))=(\prod_{i}a_{i}^{c_{i}})C(\cdot).$ Thus, lemma 3 can be applied to characterize the full set of rational functions that are invariant under such transformation, with the (CNC) reduced to ratios by picking $w_{j}:=w_{k}$ and then dividing the ratio-of-differences by $w_{k}$. ∎
@@ IV Examples
In this section, we illustrate how the methodology proposed in this paper is applicable to engineering problems comprising multiple agents competing for a limited resource. We considered two timely domains: water allocation and traffic control. We illustrate how a designer can decide the appropriate interpersonal comparability level, considering the available information and societal or political perceptions, and select an appropriate social cost function.
@@ IV-A Water allocation
In irrigation systems, the value a farmer draws from a certain amount of allocated water is difficult to quantify, especially as agricultural farmland and resulting needs are heterogeneous due to crops requiring different amounts of water per hectare and across seasons, with all of it being precipitation dependent [44, 46]. Thus, one can only conclude that farmers generally experience increased benefits with higher water allocations, though comparing it across heterogeneous farmers is not appropriate [52] (Figure 4). As stated in [53], regulators are interested in designing schemes that robustly satisfy social welfare and justice criteria despite this unmeasurable heterogeneity. These considerations, in the context of comparability, correspond to Cardinal Non Comparability of farmers’ costs.
@@
Under CNC, the appropriate social welfare function is the Nash Social Welfare. For the purpose of water allocation, farmers usually own water rights or water shares to cover the size of their land and account for the crop they are growing. In such a setting, the axiom of Anonymity (A) is not appropriate as the allocation must consider farmers’ different access rights to water. Thus, different exponents $c_{i}$ are allowed and appropriate in the SCF.
@@
In such a case, the SCF becomes $-\prod_{i}\left[\,q_{i}x^{(i)}\right]^{c_{i}}$. The solution of (2) then satisfies proportional fairness [54], which coincides with the proportional allocation
@@ IV-B Traffic Control
[…]
@@
[…]
@@
[…]
@@
[…]
@@ V Conclusions
Much of applied control work has implicitly been performed with the interests of the population in mind and based on “welfarist” principles as we laid out in this document. This work aims to provide guidance for making the foundations of the use of social cost functions explicit and, thus, to enable control theorists to express and defend their objectives based on the limits/possibilities of interpersonal comparability in various contexts and applications. In some cases, some social cost functions have to be ruled out, while many functions are permissible in other situations, and our theory thus may restrict or make explicit the range of control objectives that an engineer may pursue. The theoretical and empirical foundations that render different social cost functions preferable (e.g., the efficiency-equity tradeoff in the choice of a SCF under CFC) deserve further investigation.

### v2

# Welfare and Cost Aggregation for Multi-Agent Control: When to Choose Which Social Cost Function, and Why?

@@ I Introduction
We expect these results to be relevant in many control applications where specifications are provided in the form of social cost/welfare function [13]. The mathematical guarantees that we provide can be incorporated in the design of system-wide metrics that are now being used in multi-agent control, like Price of Anarchy [14, 15, 16, 17, 18]. They can also guide the design of optimization dynamics [19, 20, 21] and resource allocation mechanisms [22, 23]. Finally, our work responds to the growing interest in fairness in control [1, 24, 25, 26, 27, 28, 29], as many fairness notions arise naturally from careful considerations of the level of interpersonal comparability [29]. Indeed, this approach can also clarify the implicit fairness assumptions in existing game-theoretic solution concepts, such as the variational Generalized Nash Equilibrium, which has been shown to rely on a strong, and often unstated, comparability assumptions that, if violated, can lead to undesirable ’fairness’ outcomes [30].
@@
The remainder of this paper is structured as follows. In Section II, we introduce desirable axioms that will set the stage for a principled welfarist approach based on social choice theory [34, 35, 36]. In Section III, we review and extend results on how desirable axioms, together with limits/possibilities of interpersonal cost comparability, determine which social cost aggregations and operations on them are permissible. In Section IV, we illustrate the versatility of the proposed framework via examples from water distribution, transportation, and energy.
@@ II Welfarism
In this section we introduce preliminaries adopted from social choice theory that allow a rigorous aggregation of individual costs and derivation of social cost functions. We lay out the foundations of a so called “welfarist´’ approach, that requires that all the relevant information for the social decision is contained in the agents’ cost functions.
@@ II-A Preliminaries and Axiomatic Foundations
Let $\mathcal{N}=\{1,2,\dots,n\}$ be a finite set of agents and let $\mathcal{X}$ denote the set of feasible outcomes $x$. An outcome can be a specific allocation of a scarce resource or any control decision that affects the agents. Each agent $i$ attaches a (possibly negative) cost $J_{i}:\mathcal{X}\to\mathbb{R}$ to every outcome, lower numbers being preferable, with $J_{i}(\cdot)\in\mathcal{J}$, where $\mathcal{J}$ is the set of all real-valued functions on $\mathcal{X}$. We let $\mathbf{J}=(J_{1},\dots,J_{n})$ denote a profile of such costs for the entire set of agents. When formulating a resource allocation problem that depends on agents’ individual evaluations represented by $\mathbf{J}$, one needs to define a single social preference relation $\succsim_{\mathbf{J}}$ on $\mathcal{X}$, so as to capture the collective or “social” viewpoint.
@@
In control applications, it is convenient to work with real-valued measures of social cost rather than a social ordering of outcomes. This requires an additional mild continuity assumption, which postulates that if an outcome $x$ is socially strictly preferred over $y$, it should remain so under a small enough perturbation of the individual costs. As we show in Section III-B, this condition is easily verified for the setting considered in this work.
@@ II-B Welfarism
As mentioned in the introduction, several well-known examples of SCFs $C$ have been proposed to aggregate individual costs. For instance, the utilitarian approach, often attributed to Bentham [38] as later formalized by Harsanyi [39], aggregates costs via a sum: $C_{\mathrm{util}}(J_{1},\dots,J_{n})=\sum_{i=1}^{n}J_{i}$. By contrast, the Rawlsian or min-max rule, inspired by Rawls [40], focuses on the well-being of the worst-off agent: $C_{\mathrm{Rawls}}=\max_{i\in\mathcal{N}}J_{i}$. Another example is the Nash Social Welfare function [41, 42], which is defined as $C_{\mathrm{Nash}}(u_{1},\dots,u_{n})=-\prod_{i=1}^{n}(-J_{i})$. Notably, these three canonical SCFs can all be viewed as members of the family of Hölder means. Generally, a large body of literature guides the selection of a SCF based on the desirable properties reflecting notions of fairness, equity (e.g. envy-freeness and its relaxations, proportional fairness, Pigou-Dalton principle etc.), efficiency, or some trade-off between the two, reflected in measures such as maximin share (MMS) guarantee, alpha-fairness, or statistical inequality indices (e.g., the Gini coefficient, coefficient of variation, or Hoover index) [43, 5, 6, 44]. Instead, the focus of this work is to provide a “first-principle” way to make this choice by first considering the fundamental assumptions of interpersonal comparability of the agents’ costs, which are needed to properly define “efficiency” and “fairness” in the first place, and from which the appropriate aggregation rule naturally follows.
@@ III-A Comparability Levels
Let $C:\mathbb{R}^{n}\to\mathbb{R}$ be a social cost function. A vector of transformations $\{\phi_{i}\in\Phi:i\in\mathcal{N}\}$ is called an invariance transformation under $C$ if, for any cost profile $\mathbf{J}=(J_{1},\dots,J_{n})$, the transformed profile
@@ III-B Permissible Social Cost Functions
Let $\mathfrak{F}:\mathcal{J}^{n}\to\mathcal{R}$ be a Social Cost Functional satisfying (IIA) ++ (P) (unless stated otherwise). We have the following results, for each level of interpersonal comparability.
@@
with $g:\mathbb{R}^{n}\to\mathbb{R}$ homogeneous of degree 1.333gg is homogeneous of degree 1 if $g(\lambda x)=\lambda\,g(x)$ for all $\lambda>0$. For example one could choose $\gamma\max(\cdot)$ with $0\leq\gamma\leq 1$, thus reflecting a balance between efficiency and equity.
@@
$\prod_{i}\left[J_{i}(x_{0})-J_{i}(x)\right]^{c_{i}}$
with benchmark $x_{0}$
$\prod_{i}\left[-J_{i}(x)\right]^{c_{i}}$
with $b_{i}=0$
and negative costs
$\sum_{i}c_{i}\,J_{i}(x)$
$\frac{1}{n}\sum_{i}J_{i}(x)\;+\;g\left(\left[\begin{smallmatrix}J_{1}(x)-\frac{1}{n}\sum_{i}{J}(x)\\ \vdots\\ J_{n}(x)-\frac{1}{n}\sum_{i}{J}(x)\end{smallmatrix}\right]\right)$
$g$ homogeneous of degree 1
@@ III-C Admissible Operations on Social Cost Functions
One needs to be particularly careful about using SCFs for these purposes because Lemma 1 only certifies their use for selecting the best social choice. It is allowed, however, to perform other operations on the values of SCF for different outcomes, as long as the result of such operations is invariant with respect to the transformations listed in Section III-A, which guarantees that the result of such operations is meaningful. This is demonstrated in Proposition 1.
@@
Can be found in [47]. ∎
@@ IV Examples
In this section, we illustrate how the methodology proposed in this paper is applicable to engineering problems comprising multiple agents competing for a limited resource. We considered three timely domains: water allocation, traffic control, and curtailment of renewable energy. We illustrate how a designer can decide the appropriate interpersonal comparability level, considering the available information and societal or political perceptions, and select an appropriate social cost function.
@@ IV-A Water Allocation
In irrigation systems, the value a farmer draws from a certain amount of allocated water is difficult to quantify, especially as agricultural farmland and resulting needs are heterogeneous due to crops requiring different amounts of water per hectare and across seasons, with all of it being precipitation dependent [48, 50]. Thus, one can only conclude that farmers generally experience increased benefits with higher water allocations, though comparing it across heterogeneous farmers is not appropriate [56] (Figure 4). As stated in [57], regulators are interested in designing schemes that robustly satisfy social welfare and justice criteria despite this unmeasurable heterogeneity. These considerations, in the context of comparability, correspond to Cardinal Non Comparability (CNC) of farmers’ costs.
@@
Under (CNC), the appropriate social welfare function is the Nash Social Welfare. For the purpose of water allocation, farmers usually own water rights or water shares to cover the size of their land and account for the crop they are growing. In such a setting, the axiom of Anonymity (A) is not appropriate as the allocation must consider farmers’ different access rights to water. Thus, different exponents $c_{i}$ are allowed and appropriate in the SCF.
@@
In such a case, the SCF becomes $\prod_{i}\left[\,q_{i}x^{(i)}\right]^{c_{i}}$. The solution of (2) then satisfies proportional fairness [58], which coincides with the proportional allocation
@@ IV-B Traffic Control
$O_{1}$$O_{2}$$D\bullet\bullet\bullet\bullet\bullet\bullet\bullet\bullet$$40$
@@
$\bullet\bullet\bullet\bullet$
$\bullet\bullet\bullet\bullet$
@@
$\bullet\bullet\bullet\bullet$
$\bullet\bullet\bullet\bullet\bullet$
@@
$\bullet\bullet\bullet\bullet$
$\bullet\bullet\bullet\bullet$
@@
IV-C Energy Curtailment
Electrical power distribution grids host an ever-increasing amount of renewable power generation (e.g., residential solar panels). These grids have finite power transfer capacity, dictated by the physical limits of the infrastructure, which is expensive and sometimes impossible to reinforce. Therefore, access needs to be regulated: when overproduction occurs and power needs to be exported from these generators to the rest of the grid, some of the generation needs to be curtailed (up to 10% of the power generated by new installations [64]).
In mathematical terms, for a pool of $N$ generators, the grid operators need to decide the power curtailments $\{x_{i}\}_{i=1}^{N}$ for each generator $i$. In deciding that, they need to ensure that the grid operational constraints (voltage and line current limits) are satisfied when generators produce $p_{i}-x_{i}$, where $p_{i}$ is their current potential production. Such a feasibility problem leaves the grid operators to decide what is the “best” curtailment: on the one hand, the transition towards a sustainable power system calls for maximizing the total generation from renewable sources (in some cases, by law [65]). On the other hand, the energy system is a shared infrastructure, and each generator expects fair access to it. See [66] for a review of fairness in energy systems and [67, 68, 69, 70, 71, 72] for concrete proposals for fair curtailment strategies. These works propose multiple solutions but provide little guidance about which one to select, except for an a posteriori quantitative assessment of various fairness indices, which are equally hard to choose and justify.
Comparability
Different comparability assumptions are possible. We give some examples.
One could argue that each MW of curtailed power is fundamentally comparable and identical (for example, because the generators are in the same price zone and incur the same financial cost), implying $J_{i}(x)=x_{i}$ and (CUC).
Alternatively, one can acknowledge the heterogeneous nature of these small stakeholders (residential users, solar farms, etc.) and conclude that no cardinal comparability of the financial cost is possible. However, it is possible to postulate when two generators are treated equitably, and assume (OLC). For example, with $J_{i}(x)=x_{i}$, (OLC) means that two generators are treated equally if the same amount of power is curtailed, as they incur the same financial cost. If instead $J_{i}(x)=x_{i}/p_{i}$, where $p_{i}$ is the power that generator $i$ would be able to produce, then (OLC) means that generators are treated equally if the same fraction of power is curtailed, as they incur the same financial cost normalized to their earning opportunity and their investments. Finally, if $J_{i}(x)=x_{i}-p_{i}$, then (OLC) means that generators exporting (i.e., selling) the same amount of power to the grid are treated equally, as they are using the grid equally.
It is a duty of the designer of the curtailment policy to interpret the mandate they received from the stakeholders and from the local regulations. Arguably, it is easier to extract these sentiments by comparing and selecting a comparability notion than by evaluating curtailment policies or analyzing their effect once deployed.
Choice of SCF
The aforementioned comparability levels dictate which SCF is appropriate to use. (CUC) leaves no other choice than the utilitarian approach of maximizing $\sum_{i}x_{i}$ (which corresponds to OPF-total in [70]). Under (OLC), the appropriate SCF is necessarily $\max_{i}J_{i}(x)$. When $J_{i}(x)=x_{i}$, we recover the Egalitarian curtailment policy proposed in [68]. When $J_{i}(x)=x_{i}/p_{i}$, we recover the OPF-generation policy in [70] (also similar to the Proportional policy in [68]). Finally, when $J_{i}(x)=x_{i}-p_{i}$, we recover the Uniform Dynamic policy in [68] and the OPF-export policy in [70].
Implications of the choice of SCF
Different comparability assumptions justify different policies that have been proposed in the literature, but have not been adequately motivated. Explainability is crucial in this application, as the financial consequences are significant (e.g., the utilitarian approach results in a “water filling” solution where generators are completely curtailed starting from those connected further away from the main grid). Figure 6 provides some intuition on the different curtailment policies. We refer to [68] for a comparison of the numerical solutions to the resulting optimization problems on a benchmark distribution grid.
Figure 6: The effect of different curtailment policies (justified by different comparability assumptions) on five generators. Curtailing generators 4 and 5 is the most efficient way to satisfy the grid operational constraints. Figure adapted from [68].
Inequitable and opaque access to the electricity grid exacerbate existing social disparities [73], hinders investments [74], and may not meet the legal mandate of DSO as transparent and neutral players [75, 76].
Besides providing a principled motivation for different curtailment policies, the proposed approach warns against using fairness indices like those reviewed in [66, 69] (e.g., Gini index, Jain index, F-fairness, K-fairness). While they are useful to measure inequality, they may have unintended consequences when used as social cost functions, as they easily violate basic axioms like (P), i.e., Pareto optimality of the solution.
@@ V Conclusions
Much of applied control work has implicitly been performed with the interests of the population in mind and based on “welfarist” principles as we laid out in this paper. This work aims to provide guidance for making the foundations of the use of social cost functions explicit and, thus, to enable control theorists to express and defend their objectives based on the limits/possibilities of interpersonal comparability in various contexts and applications. In some cases, some social cost functions have to be ruled out, while many functions are permissible in other situations, and our theory thus may restrict or make explicit the range of control objectives that an engineer may pursue. The theoretical and empirical foundations that render different social cost functions preferable (e.g., the efficiency-equity tradeoff in the choice of a SCF under CFC) deserve further investigation.